the total value of quarters and nickels in a jar full of change is $7.70. If the nickels were quarters and the quarters were nickels, the total value of the coins would be $4.90. Find the number of nickels in the bank.
Found 2 solutions by richwmiller, ptaylor:
Answer by richwmiller(17219)

25q+5n=770
5q+25n=490
n=14 q=28
$7.00 in quarters and $.70 in nickels

Answer by ptaylor(2198)

OK. Lets deal in pennies.
Let x=number of nickels
And y=number of quarters
Now we are told the following:
5x+25y=770-------------------------eq1
and next we are told:
25x+5y=490----------------------eq2
multiply each term in eq2 by 5:
125x+25y=2450----------------------eq2a
subtract eq1 from eq2a and we get:
120x=1680 divide each term by 120
x=14----------------------------------number of nickels
substitute x=14 into eq1:
70+25y=770
25y=700
y=28----number of quarters
CK
5*14+25*28=770
70+700=770
770=770
and
25*14+5*28=490
350+140=490
490=490
Hope this helps---ptaylor